For using a computerized financial news network for 45 min during prime time and 80 min during non-prime time, a customer was charged $11.85. A second customer was charged $15.20 for using the network for 65 min of prime time and 90 min of non-prime time. Find the cost per minute for using the financial news network during prime time./
Answer by ewatrrr(24785)


Question States:
45x + 80y = $11.85
65x + 90y = $15.20

x(Prime time charge) = $0.13 and y = $0.075
45*$0.13 + 80*$0.075 = $11.85 checks